# PathM3: A Multimodal Multi-Task Multiple Instance Learning Framework for Whole Slide Image Classification and Captioning

### v1

@@ 1 Introduction
Histopathology remains the gold standard for diagnosing a wide range of cancers[17, 14]. With the rise of deep learning techniques, computational histopathology has made remarkable advances, especially in training models on gigapixel whole slide images (WSIs) from Hematoxylin and Eosin (H&E)-stained specimens [13]. Pathologists typically write diagnostic captions informed by their analysis of WSIs. These captions contribute valuable insights to the diagnostic process. The potential of utilizing such expert knowledge has increasingly attracted interest in developing deep learning models that can process both WSIs and captions to yield more comprehensive and interpretable diagnostic outcomes.
@@
Recent vision-language models have achieved significant success in nature images and texts[16, 1, 9, 10]. Among these, the query-based transformer architecture has been particularly effective due to its capability to capture fine-grained alignment between visual and textual data[1, 9]. Inspired by these developments, the trend towards multimodal learning combining image and text modalities has expanded into the medical domain[21, 25, 11, 15, 20]. Based on these prior studies, we aitm to adapt the query-based transformer to the domain of computational histopathology, with a special focus on the fusion of WSIs and WSI-level diagnostic captions. Nevertheless, integrating WSIs with WSI-level diagnostic captions presents unique challenges. First, WSIs are not suitable for direct input into deep learning models due to their immense size. Moreover, unlike natural images, which are normally independently and identically distributed, the patches extracted from WSIs exhibit redundancy and correlation, which demand specific attention or processing techniques. Second, reliable WSI diagnostic captions require specialized pathologists and are limited by privacy concerns, leading to a scarcity of such captions vital for training effective models.
@@
Multiple Instance Learning (MIL) serves as a popular solution to the first challenge. It processes the instances derived from a “bag” as inputs and predicts the bag-level label. Each WSI is considered as a “bag” and the extracted patches are regarded as instances within this bag. Considering the correlations among instances, we employ an aggregation mechanism to combine instance features into bag-level representation, which can then be used as the input for a query-based transformer for modality fusion. Existing studies mainly address the shortage of WSI-level captions by developing datasets through the use of books, articles, and web sources to compile large-scale histopathology image-caption pairs [5, 6]. However, these efforts generally yield captions limited to the patch level rather than the WSI level. In response, we propose a multi-task joint learning framework that supports both WSI classification and captioning task. This framework aims to maximize the utility of limited captions and enhance learning efficiency and predictive accuracy by leveraging multimodal data sources.
@@ 3.0.3 WSI and Caption Fusion
We utilize a query-based transformer to establish a connection between WSIs and their corresponding captions. This module takes $K$ learnable query embeddings as inputs. Initially, these queries interact with each other through self-attention layers. When textual data is available, the query further engages with textual information within the same self-attention layers. Following self-attention, the queries proceed to interact with image features via cross-attention. The output of this module is a set of $K$ integrated visual-text vectors, with each vector corresponding to a specific query embedding. After passing through $N$ blocks, these output query vectors are then fed into various specialized modules, tailored for specific tasks.
@@ 3.0.4 Multi-task Joint Learning
TieNet [21] leverages multi-task joint learning on X-ray images. We extend this approach to the domain of WSIs. Our PathM3 employs a multi-task joint learning framework by using WSIs as the sole input for inference, while taking advantage of both WSIs and captions during the training phase. This strategy allows us to harness the synergistic effects of multi-modal data, resulting in improved accuracy for both classification and captioning task. Specifically, for the classification task, we introduce both WSIs and captions into our fusion module, which outputs $K$ query embeddings that encapsulate combined visual-textual information. Subsequently, each of these output query embeddings is individually passed through a linear classifier to generate $K$ corresponding logits. Then, the classification prediction $p_{i}$ is obtained by averaging these $K$ logits across all query embeddings. To optimize the classification performance, we employ the cross-entropy function $L_{C}$, which allows us to compute the classification loss
@@
$L_{\text{overall}}=\alpha L_{C}+(1-\alpha)L_{R},$ (4)
@@ 4.3 Ablational studies
$0.503\pm 0.006$
$0.386\pm 0.002$
$0.579\pm 0.005$
$0.508\pm 0.012$
$0.388\pm 0.005$
$0.582\pm 0.011$
$0.508\pm 0.169$
$0.388\pm 0.006$
$\textbf{0.591}\pm\textbf{0.006}$
$\textbf{0.520}\pm\textbf{0.011}$
$\textbf{0.394}\pm\textbf{0.006}$
$\textbf{0.591}\pm\textbf{0.011}$
@@ 5 Conclusion
BLEU@4
Image id
Original caption
Predicted caption
0
12ffb560b9b54be8bf574c9e96804dcb
highly columnar epithelium with disordered nuclear localization proliferates showing fusion ductal construction
in the superficial epithelium, tumor tissue that invades by forming medium-sized to small, irregular ducts is observed
0.34
20a86214444542ebb703f47334620827
the superficial epithelium shows a large sheet-like shape, and some tumor tissue infiltrates with small irregular ducts
the superficial epithelium has a large amount of tumor tissue that infiltrates with small irregular ducts
0.58
76bec5ab16da4801b5a707c9c1a30542
in the superficial epithelium, tumor tissue that invades by forming medium-sized to small, irregular ducts is observed. tumor cells are columnar or cuboidal in shape
in the superficial epithelium, tumor tissue that invades by forming medium-sized to small, irregular ducts is observed
0.85
debebb808a1e45ea95656599098fea9c
tumor tissue consisting of cord-like or small, irregular glandular ducts fused and infiltrated is observed in the superficial epithelium
tumor tissue consisting of cord-like or small, irregular ducts fused and infiltrated is observed in the superficial epithelium
1
17770d23ffc8494c829b82591a73f7a5
from the superficial epithelium to the muscularis mucosae, tumor tissue consisting of medium-sized and irregular glandular ducts infiltrating is observed
from the superficial epithelium to the muscularis mucosae, tumor tissue consisting of medium-sized and irregular glandular ducts infiltrating is observed
@@
Learning rate
$41e-$ 4
Warmup learning rate
$51e-$ 5
Warmup steps
1000
Weight decay
0.05
AdamW $\beta$
$(0.9,0.999)$
Batch size
16
Image Encoder
ViT-g/14
LLM
@@
Query embedding numbers
32
Query embedding dimension
768
Hidden dimension of query-based transformer
768
Layer dimension of query-based transformer
12
Hidden dimension of correlation module
1408

### v2

# Colleyville Heritage High School, Colleyville, TX 76034, USA

@@ 1 Introduction
Histopathology remains the gold standard for diagnosing a wide range of cancers[31, 10]. With the rise of deep learning techniques, computational histopathology has made remarkable advances[25, 22], especially in training models on gigapixel whole slide images (WSIs) from Hematoxylin and Eosin (H&E)-stained specimens [27]. Pathologists typically write diagnostic captions informed by their analysis of WSIs. These captions contribute valuable insights to the diagnostic process. The potential of utilizing such expert knowledge has increasingly attracted interest in developing deep learning models that can process both WSIs and captions to yield more comprehensive and interpretable diagnostic outcomes.
@@
Recent vision-language models have achieved significant success in nature images and texts [16, 1, 11, 12, 28]. Among these, the query-based transformer architecture has been particularly effective due to its capability to capture fine-grained alignment between visual and textual data[1, 11]. Inspired by these developments, the trend towards multimodal learning combining image and text modalities has expanded into the medical domain[20, 29, 13, 15, 19]. Based on these prior studies, we aitm to adapt the query-based transformer to the domain of computational histopathology, with a special focus on the fusion of WSIs and WSI-level diagnostic captions. Nevertheless, integrating WSIs with WSI-level diagnostic captions presents unique challenges. First, WSIs are not suitable for direct input into deep learning models due to their immense size. Moreover, unlike natural images, which are normally independently and identically distributed, the patches extracted from WSIs exhibit redundancy and correlation, which demand specific attention or processing techniques. Second, reliable WSI diagnostic captions require specialized pathologists and are limited by privacy concerns, leading to a scarcity of such captions vital for training effective models.
@@
Multiple Instance Learning (MIL) serves as a popular solution to the first challenge[26]. It processes the instances derived from a “bag” as inputs and predicts the bag-level label. Each WSI is considered as a “bag” and the extracted patches are regarded as instances within this bag. Considering the correlations among instances, we employ an aggregation mechanism to combine instance features into bag-level representation, which can then be used as the input for a query-based transformer for modality fusion. Existing studies mainly address the shortage of WSI-level captions by developing datasets through the use of books, articles, and web sources to compile large-scale histopathology image-caption pairs [6, 7]. However, these efforts generally yield captions limited to the patch level rather than the WSI level. In response, we propose a multi-task joint learning framework that supports both WSI classification and captioning task. This framework aims to maximize the utility of limited captions and enhance learning efficiency and predictive accuracy by leveraging multimodal data sources.
@@ 3.0.3 WSI and Caption Fusion
We utilize a query-based transformer to establish a connection between WSIs and their corresponding captions. Similar to Flamingo[1] and Blip2[11], this module takes $K$ learnable query embeddings as inputs. Initially, these queries interact with each other through self-attention layers. When textual data is available, the query further engages with textual information within the same self-attention layers. Following self-attention, the queries proceed to interact with image features via cross-attention. The output of this module is a set of $K$ integrated visual-text vectors, with each vector corresponding to a specific query embedding. After passing through $N$ blocks, these output query vectors are then fed into various specialized modules, tailored for specific tasks. Since our query-based transformer adds cross-attention layers to BERTbase[4], similar to Blip2[11], we initialize it with the pre-trained weights from the first stage of Blip2[11].
@@ 3.0.4 Multi-task Joint Learning
TieNet[20] leverages multi-task joint learning on X-ray images. We extend this approach to the domain of WSIs. Our PathM3 employs a multi-task joint learning framework by using WSIs as the sole input for inference, while taking advantage of both WSIs and captions during the training phase. This strategy allows us to harness the synergistic effects of multi-modal data, resulting in improved accuracy for both classification and captioning task. Specifically, for the classification task, we introduce both WSIs and captions into our fusion module, which outputs $K$ query embeddings that encapsulate combined visual-textual information. Subsequently, each of these output query embeddings is individually passed through a linear classifier to generate $K$ corresponding logits. Then, the classification prediction $p_{i}$ is obtained by averaging these $K$ logits across all query embeddings. To optimize the classification performance, we employ the cross-entropy function $L_{C}$, which allows us to compute the classification loss:
@@
$L_{\text{overall}}=\alpha L_{C}+(1-\alpha)L_{G},$ (4)
@@ 4.3 Ablational studies
$0.503\pm 0.006$ $0.386\pm 0.002$ $0.579\pm 0.005$
$0.508\pm 0.012$ $0.388\pm 0.005$ $0.582\pm 0.011$
$0.508\pm 0.169$ $0.388\pm 0.006$ $\textbf{0.591}\pm\textbf{0.006}$
$\textbf{0.520}\pm\textbf{0.011}$ $\textbf{0.394}\pm\textbf{0.006}$ $\textbf{0.591}\pm\textbf{0.011}$
@@ 5 Conclusion
5.0.1 Acknowledgements.
This work was partially supported by US National Science Foundation IIS-2412195, CCF-2400785 and the Cancer Prevention and Research Institute of Texas (CPRIT) award (RP230363).
5.0.2 Disclosure of Interests.
The authors have no competing interests to declare that are relevant to the content of this article.
@@
BLEU@4 Image id Original caption Predicted caption
0 12ffb560b9b54be8bf574c9e96804dcb highly columnar epithelium with disordered nuclear localization proliferates showing fusion ductal construction in the superficial epithelium, tumor tissue that invades by forming medium-sized to small, irregular ducts is observed
0.34 20a86214444542ebb703f47334620827 the superficial epithelium shows a large sheet-like shape, and some tumor tissue infiltrates with small irregular ducts the superficial epithelium has a large amount of tumor tissue that infiltrates with small irregular ducts
0.58 76bec5ab16da4801b5a707c9c1a30542 in the superficial epithelium, tumor tissue that invades by forming medium-sized to small, irregular ducts is observed. tumor cells are columnar or cuboidal in shape in the superficial epithelium, tumor tissue that invades by forming medium-sized to small, irregular ducts is observed
0.85 debebb808a1e45ea95656599098fea9c tumor tissue consisting of cord-like or small, irregular glandular ducts fused and infiltrated is observed in the superficial epithelium tumor tissue consisting of cord-like or small, irregular ducts fused and infiltrated is observed in the superficial epithelium
1 17770d23ffc8494c829b82591a73f7a5%****␣paper.tex␣Line␣225␣**** from the superficial epithelium to the muscularis mucosae, tumor tissue consisting of medium-sized and irregular glandular ducts infiltrating is observed from the superficial epithelium to the muscularis mucosae, tumor tissue consisting of medium-sized and irregular glandular ducts infiltrating is observed
@@
Learning rate $41e-$ 4
Warmup learning rate $51e-$ 5
Warmup steps 1000
Weight decay 0.05
AdamW $\beta$ $(0.9,0.999)$
Batch size 16
Image Encoder ViT-g/14
@@
Query embedding numbers 32
Query embedding dimension 768
Hidden dimension of query-based transformer 768
Layer dimension of query-based transformer 12
Hidden dimension of correlation module 1408